Determine the slope of the line that passes through each pair of points.
(-2,2) (3,-3)
(-5,4) (-1,11)
(-2,-8) (1,4)
(18,-4) (6,-10)
(3,4) (4,6)
(3,6) (4,4)
(4,-5) (11,-1)
(2,7) (-5,5)

Answer by blwinbbbles(106)

You will use the point slope equation:
Which is Y2-Y1/X2-X1
I will do the first one to show you how.
(-2,2) (3,-3)
So:
-3 - 2/3-(-2) =
-5/5 =
-1 so the slope for that equation is -1
You can now try the rest of the cordinates